Please show work and any calculator functions if available. Answer= . 0.0116

The annual amounts of snowfall in a Rocky mountain pass are normally distributed with a mean of 49.2 inches and a standard deviation of 11.5 inches. What is the probability that the mean annual amount of snowfall during 9 randomly selected years will be less than 40.5 inches? (Central Limit Theorem)
Answer #1

The following information has been provided:

μ=49.2, σ=11.5

for sampling distribution for n=9

xbar ~ μ=49.2, σ=11.5/9

μ=49.2, σ=3.8333

Z=(X−μ​)/σ

Pr(Xbar ≤ 40.5) =Pr(Z≤−2.2696)=0.0116